# SpectralEarth: Training Hyperspectral Foundation Models at Scale

### v1

@@ 3.1 SpectralEarth
The EnMAP data was sourced from the DLR GeoPortal. A total of 11,636 tiles were retrieved via the portal’s graphical interface through a significant manual effort. All tiles were carefully selected based on human visual inspection. Cloud coverage was kept below $\sim$10%. The selection process encompassed the entire EnMAP archive from April 2022 to April 2024. All acquired tiles underwent radiometric, geometric, and atmospheric corrections, resulting in Level-2A products.
Data Pre-processing.
@@
The EnMAP tiles have been split into geospatial patches—each of size 128x128 pixels where an individual pixel includes 224 bands. Bands dominated by water absorption were excluded due to the frequent presence of missing values (no-data). Removing those badns is a common practice in hyperspectral dataset creation [19, 60]. As a result, the total number of bands per image reduced to 202.
Temporal Views Extraction.
@@
To exploit all available tiles, we utilized overlaps between EnMAP data acquisitions to generate time series of EnMAP patches [^1] EnMAP records hyperspectral imagery based on a schedule generated from user requests [37], i.e. the identification of spatially aligned patches for multiple timestamps is a valuable asset our data curation provides.. At the same time SpectralEarth data patches for a fixed geo-location never overlap. This process required (a) identification of tiles that overlap spatially to (b) extract patches from the intersection of those tiles. Time series include valuable information on landscape evolution and provide characteristic signals for seasonal variation—an asset for contrastive learning [46, 65]. Algorithm 1 outlines the steps taken to generate temporal positives by high-level pseudo-code.
@@
6:         combinations
[…]
8:         for tile_subset in combinations do
9:             intersection $\leftarrow$ Intersection(tile_subset)
10:             patches $\leftarrow$ Patchify(intersection)
11:             Update(R_tree, patches)
@@
13:     end for
14:end procedure
15:function BuildCombinations(tile, overlap_graph)
16:    combinations $\leftarrow$ GetEdges(tile, overlap_graph)22footnotemark: 2
17:    for subset size n in $[3,4,\dots]$ do
18:         get n-tuples from (n-1)-tuples in combinations
19:         compute intersections of all n-tuples
20:         keep largest n-tuples by area
21:         if no valid n-tuple found then
22:             break
23:         end if
24:         add n-tuples to combinations
25:    end for
26:    return combinations
27:end function
As a result of Algorithm 1 we identified a few long time series in the EnMAP archive. Those involving more than 25 intersecting tiles. Calculating intersections of all subsets was computationally intractable due to combinatorial explosion. To address this, we implemented several heuristics to optimize the computation:
@@
Following this approach, we extracted 73,307 non-overlapping locations with multiple timestamps. For areas with a single timestamp we patchified avoiding any spatial overlap. Correspondingly, in SpectralEarth we extracted 415,153 locations with a spatial extent of 128x128 pixels, totaling 538,974 patches. Figure 6 illustrates the distribution of timestamp counts in SpectralEarth. Over the two-year’s period, 17.5% of locations were covered by more than one timestamp in the EnMAP archive—with the majority having only two timestamps. This highlights the challenge of collecting extensive hyperspectral time series data with broad geographical coverage today. Missions such as ESA’s CHIME satellite constellation [53] will eventually resolve the issue.
Figure 6: Histogram representing the number of timestamps (log-scale) for geo-locations covered by SpectralEarth. The majority of locations include one timestamp, while the remaining ones mostly consist of pairs and triplets.
The spatial coverage of SpectralEarth showcases the dataset’s global diversity, cf. Figure 3. Notably, the data distribution varies across geo-locations. EnMAP is a mission with scheduled acquisition depending on areas of interest.
3.2 Downstream Tasks
@@
To evaluate the models pretrained on SpectralEarth, we assembled four downstream datasets for benchmarking. Each benchmark involves a subset of EnMAP or DESIS images, aligned with specific geospatial products with focus on specific aspects of land cover and agricultural analysis. While the labels in these datasets carry some inherent uncertainty, they are sufficiently accurate for a proof-of-concept evaluation of model performance. Similar strategies have previously been employed to create remote sensing datasets [56, 57, 34].
EnMAP-CORINE: Land Cover Classification.
@@
This dataset pairs EnMAP imagery with the CORINE land cover database for Europe. The CORINE land cover map has a 100m spatial resolution, which is coarse-grained relative to EnMAP’s 30m resolution. Our SpectralEarth image patches correspond to approximately 38x38 CORINE land cover pixels. Consequently, we create a multi-label classification benchmark where each SpectralEarth patch is annotated with the set of CORINE classes covered. The resulting subset, called EnMAP-CORINE, includes 11,000 patches, each multi-labeled with 19 distinct classes. These classes have been aggregated from the 44 categories defined in CORINE following the taxonomy proposed by BigEarthNet-MM [58]. Example patches from four diverse scenes are depicted in Figure  4.
EnMAP-CDL: Crop Type Segmentation.
@@
This dataset aligns EnMAP images with the Cropland Data Layer (CDL) [8] product for the United States. The CDL map is published annually by the USDA’s National Agricultural Statistics Service since 2008. It targets agricultural crop classification. In addition to classes for cultivated land such as Soybeans and Winter Wheat, CDL encompasses non-cultivated areas like Grassland and Shrubland. To minimize the impact of seasonal variations on crop classification, the dataset utilizes SpectralEarth patches from the summers of 2022 and 2023. Given the strong imbalance among CDL classes, we resampled the class distribution to set up a more balanced representation. The resulting dataset comprises 2,000 patches with a total of 20 classes.
EnMAP-NLCD: Land Cover Segmentation.
@@
This downstream dataset leverages the National Land Cover Database (NLCD) [17] from the U.S. Geological Survey (USGS) to provide pixel-level land cover annotations. The NLCD map, a comprehensive land cover product, has been published every 2–3 years since 2001, with data available back to 2019. The resulting dataset contains 15,000 patches, each annotated with pixel-wise land cover class labels from one out of the 15 land cover classes.
DESIS-CDL: Crop Type Segmentation.
@@
The DESIS-CDL dataset contains images captured for the summers of 2021, 2022 and 2023 from the DESIS (DLR Earth Sensing Imaging Spectrometer Mission) instrument, matched with the corresponding CDL mask for crop segmentation. DESIS images are generated at 30m spatial resolution with 235 bands covering wavelengths from 400 nanometers through a micron. The resulting dataset comprises 1,000 images with 19 classes.
4 Hyperspectral Foundation Models
@@
4.1 Pre-training Algorithms
Hyperspectral data has very different characteristics compared to natural images. Therefore, traditional SSL leaderboards in computer vision may not directly translate to the hyperspectral domain. To address this gap, we explore a broad range of SSL methods to establish a set of baseline pretrained models for hyperspectral imagery. We use MoCo-V2 [26] and DINO [12] for their effectiveness in frozen encoder performance [51] and their compatibility with both CNNs and ViTs. Furthermore, we include MAE [27] for its strong performance in fine-tuning scenarios. Together, these methods form a representative set of algorithms to investigate SSL in the hyperspectral domain.
4.2 Network Architectures
@@
  • Computational Efficiency: 3D convolution or tokenizing the image in the spectral dimension is computationally expensive for large models. We seek for adaptation of classical architectures with minimal overhead.
@@
To meet these requirements, we follow a pragmatic approach and implement a simple modular design, integrating 1D convolutional layers as a Spectral Adapter at the onset of the standard vision backbones, as depicted in Figure 7. These layers perform convolutions across the spectral dimension, generating feature maps for the core backbone. Specifically:
  • Spectral ResNet: We replace the stem layers with three 1D convolutional layers and pooling, allowing the network to process spectral information effectively before the residual backbone. By removing the original stem layers in the ResNet architecture, we avoid the 4x downscaling factor of the image to better capture fine-grained details.
@@
  • Spectral ViT: The 1D convolutions and pooling are placed before the patch embedding layer, ensuring that the transformer receives fixed-size spectral features as input. We use 4x4 patches instead of the typical 16x16 patches, which helps maintain fine spatial details and better preserves spectral information at the patch projection layer.
This simple strategy results in an architecture that can be used on various hyperspectral sensors while leveraging a lightweight adapter for spectral feature extraction. Similar designs have been used effectively for spectral-spatial classifiers in the HSI literature [1]. Our approach extends this idea to foundation models, enabling the extraction of both spatial and spectral features while avoiding the extra computational cost incurred by 3D convolutions or spatial-spectral tokenization in ViTs.
5 Experimental Setup
5.1 Self-Supervised Pre-training
@@
We pre-train a spectral ResNet-50 (Spec. RN50) and a spectral ViT-S (Spec. ViT-S) with MoCo-V2, DINO, and MAE. For larger models, Spec. ViT-Base(B)/Large(L)/Huge(H)/Giant(g), we restrict the experiments to MAE given the high computational cost. For ViT-g, we follow the implementation from [51]. We use the augmentations from SimCLR [14] without color-jittering and gray-scale for DINO and MoCo-V2 to avoid distorting spectral information. When available, we use the temporal views as positive pairs for MoCo-V2 and DINO. For MAE, we mask $90\%$ of the tokens. We pretrain over 100 epochs for MoCo-V2 and DINO, and 200 epochs for MAE with a batch size of 256 images. We use a StepLR scheduler for MoCo-V2, and CosineAnnealingLR for DINO and MAE.
5.2 Downstream Tasks
Multi-label classification on EnMAP-CORINE.

### v2

@@ III-A SpectralEarth
The EnMAP data was sourced from the DLR GeoPortal. A total of 11,636 tiles were retrieved via the portal’s graphical interface through a significant manual effort. All tiles were carefully selected based on human visual inspection. Cloud coverage was kept below $\sim$10%. The selection process encompassed the entire EnMAP archive from April 27, 2022 to April 24, 2024. All acquired tiles underwent radiometric, geometric, and atmospheric corrections using the L2A processor from the EnMAP mission [71].
Data Pre-processing
@@
The EnMAP tiles have been split into geospatial patches—each of size $\times128$ pixels where an individual pixel includes 224 bands. Bands dominated by water absorption, specifically [127–141] and [161–167], were excluded due to the frequent presence of missing values (no-data). Removing those bands is a common practice in hyperspectral dataset creation [14, 32]. As a result, the total number of bands per image was reduced to 202.
Temporal Views Extraction
@@
To exploit all available tiles, we utilized overlaps between EnMAP data acquisitions to generate time-series of EnMAP patches [^2] EnMAP records hyperspectral imagery based on a schedule generated from user requests [7], i.e. the identification of spatially aligned patches for multiple timestamps is a valuable asset our data curation provides.. At the same time, SpectralEarth patches for fixed geo-locations never overlap. This process required (a) the identification of tiles that overlap spatially and (b) extracting patches from the intersection of those tiles. Time series include valuable information on landscape evolution and provide characteristic signals for seasonal variation—an asset for contrastive learning [47, 48]. Algorithm 1 outlines the steps taken to generate temporal positives by high-level pseudo-code. An overlap graph is first constructed to identify spatial intersections between tiles. For each tile, the algorithm forms candidate subsets of overlapping tiles, computes their joint intersection, and extracts patches of fixed size from these regions. We use an R-tree to efficiently track and store patch locations and avoid overlaps.
@@
6:    combs $\leftarrow$ BuildCombinations(tile, overlap_graph)
7:    for tile_subset in combs do
8:     intersection $\leftarrow$ Intersection(tile_subset)
9:     patches $\leftarrow$ Patchify(intersection)
10:     INSERT(R_tree, patches)
11:    end for
@@
13:end procedure
14:function BuildCombinations(tile, overlap_graph)
15:  combinations $\leftarrow$ GetEdges(tile, overlap_graph)33footnotemark: 3
16:  for subset size n in $3,4,\dots]$ do
17:    get n-tuples from (n-1)-tuples in combinations
18:    compute intersections of all n-tuples
19:    keep largest n-tuples by area
20:    if no valid n-tuple found then
21:     break
22:    end if
23:    add n-tuples to combinations
24:  end for
25:  return combinations
26:end function
As a result of Algorithm 1, we identified a few long time series in the EnMAP archive. For those involving more than 25 intersecting tiles, calculating intersections of all subsets was computationally intractable due to combinatorial explosion. To address this, we implemented several heuristics to optimize the computation:
@@
(a) Number of timestamps per geo-location (log scale).
(b) Gap between consecutive timestamps.
(c) Monthly distribution aggregated over 2 years.
Figure 5: Temporal distribution of SpectralEarth: (a) number of timestamps per location, (b) time gaps between acquisitions, and (c) monthly acquisition frequency.
Following this approach, we extracted 73,307 non-overlapping locations with multiple timestamps. For areas with a single timestamp, we patchified avoiding any spatial overlap. Correspondingly, in SpectralEarth, we extracted 415,153 locations with a spatial extent of $\times128$ pixels, totaling 538,974 patches. The spatial coverage of SpectralEarth showcases the dataset’s global diversity in land-cover types (Figure 2) and geographical distribution (Figure 3). Notably, the data distribution varies across geo-locations, reflecting the flight request-based archive composition of EnMAP.
@@
We analyze the temporal coverage of our dataset in Figure 5. Figure 5(a) illustrates the distribution of timestamp counts in SpectralEarth. Over the two-year period, 17.5% of locations were covered by more than one timestamp in the EnMAP archive—with the majority having only two timestamps. This highlights the challenge of collecting extensive hyperspectral time series data with broad geographical coverage today. Missions such as ESA’s CHIME satellite constellation [10] will eventually resolve the issue. Figure 5(b) shows the histogram of time gaps between consecutive acquisitions, restricted to samples with at least two timestamps. We observe that the distribution is skewed towards smaller values. Notably, the highest peak is 4 days, matching EnMAP’s revisit period. The median gap between two consecutive acquisitions across all samples is 27 days.
We also analyze the seasonal distribution of EnMAP acquisitions, as shown in Figure 5(c). SpectralEarth includes data from all months, with a relatively balanced distribution. However, certain disparities can be observed, e.g., January is underrepresented while July, August, and March are over-represented. This can be attributed to several factors:
  • Increased cloud cover during winter months, which inherently limits the availability of clear imagery;
  • A satellite outage in December-January 2023 and December-January 2024, which resulted in a temporary interruption of data acquisition;
  • The task scheduling nature of the EnMAP mission. As EnMAP is operationally tasked based on specific use cases, the data distribution inherently reflects real-world demands of satellite imagery, leading to a natural variation in seasonal coverage.
(a) EnMAP-CORINE
(b) EnMAP-NLCD
(c) EnMAP-Treemap
(d) EnMAP-CDL
(e) EnMAP-EuroCrops
(f) EnMAP-BDForet
(g) EnMAP-BNETD
(h) DESIS-CDL
(i) EO1-CDL
Figure 6: Class distributions for the downstream datasets introduced in SpectralEarth.
III-B Downstream Tasks
@@
To evaluate the models pretrained on SpectralEarth, we assembled nine downstream datasets for benchmarking. Each benchmark involves a subset of EnMAP/DESIS/EO-1 images aligned with specific geospatial products focusing on different aspects of land cover, agricultural, and forest analysis. While the labels in these datasets carry some inherent uncertainty, they are sufficiently accurate for model evaluation. Similar strategies have previously been employed to create remote sensing datasets and geospatial foundation model benchmarking [72, 72, 73]. An overview of the SpectralEarth downstream tasks is provided in Figure 4.
EnMAP-CORINE - Land Cover Classification
@@
This dataset pairs EnMAP imagery with the CORINE land cover database for Europe. The CORINE land cover map has a 100m spatial resolution, which is coarse-grained relative to EnMAP’s 30m resolution. The CORINE product reports an overall thematic accuracy exceeding 85%. Our SpectralEarth image patches correspond to approximately 38x38 CORINE land cover pixels. Consequently, we create a multi-label classification benchmark where each SpectralEarth patch is annotated with the set of CORINE classes covered. The resulting subset, called EnMAP-CORINE, includes 11,000 patches, each multi-labeled with 19 distinct classes. These classes have been aggregated from the 44 categories defined in CORINE following the taxonomy proposed by BigEarthNet-MM [74]. The EnMAP-CORINE dataset exhibits a diverse range of land cover types. As Figure 6(a) demonstrates, some classes such as ’Arable Land’ and ’Complex cultivation patterns’ are more frequent, while others like ’Coastal/Inland wetlands’ and ’Beaches, dunes, sand’ are comparatively rare. This class imbalance reflects the distribution of natural land cover in Europe.
EnMAP-CDL - Crop Type Segmentation
@@
This dataset aligns EnMAP images with the Cropland Data Layer (CDL) [75] product for the United States. The CDL map is published annually by the USDA’s National Agricultural Statistics Service since 2008. It targets agricultural crop classification. In addition to cultivated crops such as Corn and Soybeans, CDL encompasses non-cultivated areas like Grassland and Shrubland. According to the USDA metadata, classification accuracy is generally between 85% and 95% for major crop-specific classes. To minimize the impact of seasonal variations on crop classification, the dataset utilizes SpectralEarth patches from the summers of 2022 and 2023. We mask out all non-agricultural classes, as well as winter crops (e.g., Winter Wheat), and crops that are likely to be harvested before the end of summer (e.g., Oats). Given the strong imbalance among CDL classes, we resampled the class distribution for a more balanced representation. The resulting dataset comprises 1,600 patches with a total of 14 classes. The class distribution of the EnMAP-CDL dataset is shown in Figure 6(d). Overall, the dataset exhibits a reasonable level of class balance among the different crop types, albeit some over-represented classes such as Corn.
EnMAP-NLCD - Land Cover Segmentation
@@
This downstream dataset leverages the National Land Cover Database (NLCD) [76] from the U.S. Geological Survey (USGS) to provide pixel-level land cover annotations. The NLCD map, a comprehensive land cover product, has been published every 2–3 years since 2001, with data available back to 2019. The NLCD 2019 product reports an overall accuracy of approximately 91%. We pair EnMAP patches over the US to the NLCD map, and manually filter out cloudy and snowy images. The resulting dataset contains 13,500 patches, each annotated with pixel-wise land cover class labels from one out of the 15 land cover classes. As depicted in Figure 6(b), the class distribution is well-balanced.
EnMAP-TreeMap - Tree Species Segmentation
The EnMAP-TreeMap dataset maps EnMAP imagery from 2022 and 2023 to TreeMap [77] product, a 30 m resolution map of forest attributes across the continental United States. We use the forest-type layer, which encodes the dominant species group based on live tree stocking. To mitigate the temporal mismatch between TreeMap (2016) and EnMAP acquisitions, we mask regions affected by deforestation using the Hansen Global Forest Change dataset [78]. The resulting dataset includes 10,000 patches covering 24 tree species, enabling fine-grained classification of forest types. The class distribution is depicted in Figure 6(c).
EnMAP-BDForet - Tree Species Segmentation
The EnMAP-BDForet dataset combines EnMAP hyperspectral imagery acquired between 2022 and 2024 with the BDForet V2 dataset from the Institut Géographique National (IGN) France  [^4] https://inventaire-forestier.ign.fr/spip.php?article646. We project the dominant tree species attribute onto the EnMAP grid to generate the labels. To mitigate land cover changes, we filter out areas affected by deforestation using the latest IGN forest mask from 2023. The resulting dataset comprises 2,550 patches with 12 tree species. The class distribution is shown in Figure 6(f).
EnMAP-EuroCrops - Crop Type Segmentation
The EnMAP-EuroCrops dataset aligns EnMAP imagery with crop type labels derived from the EuroCrops dataset [79], which is sourced from national agricultural inventories across European countries. This dataset focuses on crop-type classification for four countries—France, Germany (Brandenburg), Czechia, and Spain—for the year 2023, with additional coverage for France in 2022. Similar to CDL, we exclude crops which are likely to be harvested before our EnMAP acquisition. The resulting dataset consists of 1,800 patches with 15 crop-type classes. Approximately 68% of the patches are located in France, 27% in Spain, and the remaining 5% in Germany and Czechia. The class distribution is illustrated in Figure 6(e).
EnMAP-BNETD - Land Cover Segmentation
To explore the geographical generalizability of our models beyond pre-training regions, we introduce the EnMAP-BNETD dataset, covering land cover classification in Ivory Coast. This dataset uses the BNETD 2020 Land Cover Map  [^5] https://africageoportal.maps.arcgis.com/apps/webappviewer/index.html?id=88c2493e722546c09c2a0a8b394c4454, produced by the BNETD-CIGN with support from the European Union, and validated through extensive field campaigns in 2022 and 2023. The reported overall accuracy is 91%. We pair EnMAP hyperspectral imagery with land cover labels from this product, resulting in a dataset of 2,100 patches with 10 classes. The class distribution is shown in Figure 6(g).
DESIS-CDL - Crop Type Segmentation
@@
The DESIS-CDL dataset contains images captured for the summers of 2018-2023 from the DESIS (DLR Earth Sensing Imaging Spectrometer Mission) instrument, matched with the corresponding CDL mask for crop segmentation. DESIS images are generated at 30m spatial resolution with 235 bands covering wavelengths from 400 nanometers through a micron. The resulting dataset comprises 1,000 images with 14 classes. Due to the limited availability of DESIS tiles, it was not possible to pair this dataset with EnMAP-CDL. Therefore, the class distribution (see Figure 6(h)) and geographical coverage differ.
EO1-CDL - Crop Type Segmentation
The EO1-CDL dataset includes EO-1 Hyperion hyperspectral data paired with crop-type labels from the CDL. EO-1 Hyperion provides 220 spectral bands covering 0.357 to 2.576 micrometers, out of which we use 198. The dataset includes Hyperion scenes from 2002 to 2016, focusing on agricultural regions over the US. Due to the limited availability of Hyperion imagery in Google Earth Engine, the dataset consists of 550 patches distributed across 14 crop classes.
IV Models
@@
IV-A Pre-training Algorithms
Hyperspectral data has very different characteristics compared to natural images. Therefore, traditional SSL leaderboards in computer vision may not directly translate to the hyperspectral domain. Therefore, we explore a broad range of SSL methods to establish a set of baseline pretrained models for hyperspectral imagery. We use MoCo-V2 [19] and DINO [21] for their effectiveness in frozen encoder performance [11] and their compatibility with both CNNs and ViTs. Furthermore, we include MAE [22] for its strong performance in fine-tuning scenarios. Together, these methods form a representative set of algorithms to investigate SSL in the hyperspectral domain.
IV-B Network Architectures
@@
  • Computational Efficiency: 3D convolution or spectral-spatial tokenization becomes computationally expensive for large patches, with many spectral bands, and deep architectures. In particular, self-attention scales quadratically with the number of tokens [83], and 3D convolutions scale linearly with the number of spectral bands [84]. We therefore seek lightweight adaptation of classical architectures with minimal overhead.
@@
To meet these requirements, we follow a pragmatic approach and implement a simple modular design, integrating 1D convolutional layers as a Spectral Adapter at the onset of the standard vision backbones, as depicted in Figure 7. These layers perform convolutions across the spectral dimension, generating feature maps for the core backbone. The Spectral Adapter consists of three (1D Conv + BN + ReLu) layers. The kernel sizes for the 1D convolutions are 7, 7, and 5, with strides of 5, 5, and 3, respectively. The resulting feature maps have 128 channels. To accommodate different sensors and feed the input into standard 2D backbones, we use a global pooling layer to aggregate the remaining spectral dimensions (if any). Specifically:
[…]
@@
  • Spectral ViT: The spectral adapter is placed before the patch embedding layer, ensuring the transformer receives fixed-size spectral features as input. We use $\times4$ patches instead of the typical $\times16$ patches, which helps maintain fine spatial details and better preserves spectral information at the patch projection layer. Like the spectral ResNet model, we set the number of input channels in the projection layer to 128.
Although these simplifications may not capture all spectral and spatial interactions as explicitly as 3D convolutions or spectral-spatial tokenization, they offer a practical trade-off that scales better to large datasets and diverse sensors. Similar designs have been used effectively in HSI classification tasks [85, 13]. Our approach adapts this idea to large-scale SSL, providing a simple scalable solution for hyperspectral representation learning across sensors.
Figure 8: Schematic overview of the pre-training/fine-tuning pipeline. An encoder is pre-trained on SpectralEarth, and adapted to the several downstream tasks with task-specific heads.
V Experimental Setup
This section introduces the details of pre-training and evaluation protocols. A schematic overview of the pipeline is provided in Figure 8.
V-A Self-Supervised Pre-training
@@
We pre-train a spectral ResNet-50 (Spec. RN50) and a spectral ViT-S (Spec. ViT-S) with MoCo-V2, DINO, and MAE. For larger models, Spec. ViT-Base(B)/Large(L)/Huge(H)/Giant(g), we restrict the experiments to MAE, given the high computational cost. For ViT-g, we follow the implementation from [11]. We use the augmentations from SimCLR [38] without color-jittering and gray-scale for DINO and MoCo-V2 to avoid distorting spectral information. When available, we use the temporal views as positive pairs for MoCo-V2 and DINO. For MAE, we mask $90\%$ of the tokens. We pretrain over 100 epochs for MoCo-V2 and DINO, and 200 epochs for MAE with a batch size of 256 images. We use a StepLR scheduler for MoCo-V2 and CosineAnnealingLR for DINO and MAE.
V-B Downstream Tasks
Multi-label classification